### v1

# Improving Autoregressive Visual Generation with Cluster-Oriented Token Prediction

Abstract
Employing LLMs for visual generation has recently become a research focus. However, the existing methods primarily transfer the LLM architecture to visual generation but rarely investigate the fundamental differences between language and vision. This oversight may lead to suboptimal utilization of visual generation capabilities within the LLM framework. In this paper, we explore the characteristics of visual embedding space under the LLM framework and discover that the correlation between visual embeddings can help achieve more stable and robust generation results. We present IAR, an Improved AutoRegressive Visual Generation Method that enhances the training efficiency and generation quality of LLM-based visual generation models. Firstly, we propose a Codebook Rearrangement strategy that uses balanced k-means clustering algorithm to rearrange the visual codebook into clusters, ensuring high similarity among visual features within each cluster. Leveraging the rearranged codebook, we propose a Cluster-oriented Cross-entropy Loss that guides the model to correctly predict the cluster where the target token is located. This approach ensures that even if the model predicts the wrong token index, there is a high probability the predicted token is located in the correct cluster, which significantly enhances the generation quality and robustness. Extensive experiments demonstrate that our IAR consistently enhances the model training efficiency and performance from 100M to 1.4B, reducing the training time by half while achieving the same FID. Additionally, IAR can be applied to various LLM-based visual generation models and adheres to the scaling law, providing a promising direction for future research in LLM-based visual generation.
1 Introduction
Figure 1: When an autoregressive model predicts a wrong token, the previous methods [38, 40] may predict an irrelevant token that causes artifacts. Our IAR alleviates this issue by ensuring a high probability of the predicted token located in the correct cluster.
With the development of generative models [35, 16, 20], a large number of outstanding visual generation methods have emerged, achieving considerable success in image and video generation [35, 4, 17]. Recently, researchers have begun exploring the integration of images and text to achieve a unified multi-modal model for image and text generation and understanding [45, 39]. Consequently, aligning image generation models with large language models (LLMs) [34] has become a key research focus. Recent studies have made progress by quantizing images and employing autoregressive or mask-prediction methods to predict discrete image tokens [38, 40, 7, 3], laying the research foundation for a unified world model in the future.
Unlike diffusion [20] or GAN [16] models that model image distribution in continuous space, autoregressive or masked image modeling (MIM) methods [38, 40, 7, 3] first convert images into discrete-valued tokens by image tokenizers, and then predict image tokens by autoregressive or MIM approach. Specifically, autoregressive methods use the “next-token prediction” paradigm of GPT [34], while MIM methods adopt a similar training strategy to the mask prediction in BERT [12]. These methods draw inspiration from natural language models and transfer these techniques to image generation models, but rarely explore the fundamental differences between images and natural language.
Due to the inherently discrete nature of words (each word comes from a finite vocabulary), different words can be directly mapped to different indices through a lookup table, thereby generating text by predicting the corresponding index of the target text. The standard language modeling objective $p(x_{t}|x_{1},\dots,x_{t-1})$ (each $x_{t}$ is an index) then can be interpreted as a classification task and accurately modeled by LLMs. However, images are continuous-valued, and if the same objective is applied to image generation with $x_{t}$ as a real number, it is difficult to accurately model this probability density. Therefore, image tokenizers such as VQGAN [14] are used to convert continuous images into discrete-valued tokens, and then the model predicts the image token index. Afterwards, the corresponding image embedding is retrieved from the codebook and decoded into an image. A fundamental difference between image and text is that, in text generation, the predicted text index can be directly mapped to the corresponding word through a lookup table; whereas LLM-based image generation essentially requires the image embedding that corresponds to the index, and then decodes it into an image. I.e., text only requires the index, while image requires the embedding corresponding to the index.
Different from the token index which is independent, i.e., the nearby indices have no correlation in word semantics, the image embeddings are located in a continuous feature space, where similar embeddings may correspond to similar image contents. This paper first investigates the impact of image embedding correlations on the generated results, and finds that similar embeddings convey similar information in the image space: When the embeddings of certain image patches are replaced by similar image embeddings, the decoded images are nearly identical. This insight inspires us to leverage the similarity of image embeddings to improve the existing LLM-based image generation techniques.
To capitalize on the correlation among image embeddings, we propose IAR, an Improved AutoRegressive Visual Generation Method, which can enhance the training efficiency and generation quality of the LLM-based visual generation model. Firstly, we propose a Codebook Rearrangement strategy that uses balanced K-means clustering algorithm to rearrange the embeddings in the codebook into clusters of equal sizes, where the embeddings in each cluster share high similarities and can be decoded into images with similar contents. Secondly, we observe that if a token index is wrongly predicted as a similar embedding’s index, the original token-oriented cross-entropy loss used in LLM will penalize this case, but the decoded image is actually not much different from the target image. To tolerate this case, based on the rearranged codebook, we further introduce a Cluster-oriented Cross-entropy Loss, which guides the model to predict the correct cluster that contains the target token, thereby providing the model with a broader perspective and no longer confined to a single target token. Combined with the original token-oriented cross-entropy loss, our model can consider both the target cluster and the target token. Even if the model predicts the wrong token index, with our rearranged codebook and cluster-oriented loss, there is a high probability that the token is located in the target cluster, which ensures a high similarity between the output image and the target one, effectively improving the generation quality and robustness (Fig. 1).
We develop our model based on LlamaGen [38] and compare with existing visual generation methods based on GAN, Diffusion, Autoregressive (AR), and MIM models. Extensive experiments demonstrate that, across different parameter scales of LlamaGen (ranging from 100M to 1.4B), our method consistently enhances the model’s training efficiency and performance. For the same FID, our approach reduces LlamaGen’s training time by half, and under the same number of training epochs, it effectively improves the model’s generation quality. Our method can be applied to various LLM-based visual generation models with almost no additional training costs, and adheres to the scaling law, providing a promising direction for improving future LLM-based visual generation models.
Our main contributions can be summarized as four-fold:
  • We propose IAR, an Improved Autoregressive Visual Generation method, which leverages the correlation of visual embeddings to enhance the training efficiency and performance of LLM-based visual generation models.
  • We introduce a Codebook Rearrangement strategy that uses balanced K-means clustering algorithm to rearrange the image codebook, ensuring high similarity across image embeddings within each cluster.
  • We propose a Cluster-oriented Cross-entropy Loss that relaxes the original token-oriented cross-entropy, ensuring that even if the model predicts the wrong token index, there is still a high probability that the token is in the correct cluster, thereby generating high-quality images.
  • Extensive experiments demonstrate the effectiveness of our method. Additionally, our approach can be applied to various LLM-based visual generation models, adheres to the scaling law, and provides a robust improvement direction for future LLM-based visual generation models.
2 Related Works
Figure 2: (a) The MSE and LPIPS between the source image and the decoded image with different code distances. (b) The visualization of the images decoded from different code distances. When the code distance is within a certain range (e.g., smaller than 12), the decoded image looks nearly identical to the source image. We further make use of this property to improve the LLM-based visual generation model.
2.1 Large language models
Large language models (LLMs) [34, 41, 12] have achieved significant success in natural language processing (NLP). The encoder-only structure [12, 10, 28], exemplified by BERT [12], randomly masks portions of the text and predicts the masked tokens based on the unmasked ones. With the introduction of the GPT series [34, 33, 6, 1], the decoder-only architecture gained popularity, demonstrating remarkable capabilities in language processing through next-token prediction. Additionally, open-source models such as LLaMA [30] and PaLM [9] have further advanced the NLP field. The recent success of LLMs in natural language processing has also inspired researchers to apply these models to visual generation, revealing significant potential in multi-modal generation tasks.
2.2 Visual Generation Models
Continuous-valued Visual Generation. Generative Adversarial Networks (GANs) [16, 25, 14, 43], consisting of a generator and discriminator, employ adversarial training to learn the target distribution. Diffusion Models [20, 35, 15, 13, 21, 31] iteratively add and remove noise to approximate the real data distribution, using a Markov process for the forward path and learning the reverse denoising process. These models have achieved significant success in image generation and have been extended to video generation [17, 4, 42], enhancing video quality. However, both GANs and Diffusion Models, which model continuous distributions, face challenges when integrating with discretely modeled LLMs. Thus, exploring discrete modeling methods for image generation is necessary.
Discrete-valued Visual Generation. Discrete-space visual generation models are mainly divided into autoregressive [38, 14, 40] and masked image modeling (MIM) [7, 8, 3] models, both requiring image quantization before generation. Autoregressive models follow the next-token generation approach of GPT [34, 1], predicting subsequent tokens one by one based on previous image tokens. LLaMAGEN [38], a representative work, uses LLAMA [41] to model the image generation process. VAR [40] further enhances generation capability by proposing a next-scale image generation process. MIM methods [38, 40, 7, 3], inspired by BERT [12], randomly mask parts of the image tokens and predict the masked tokens using the unmasked ones. These methods directly draw inspiration from LLMs and rarely explore the fundamental differences between images and language, which may lead to suboptimal utilization of vision generation capabilities within the LLM framework. In contrast, our model makes full use of the image property, which effectively improves the training efficiency and model performance in LLM-based image generation.
3 Method
Figure 3: Model framework: 1) Codebook Rearrangement: we first use a balanced K-means clustering method to rearrange the codebook, which divides the codebook into $n$ clusters, with the image codes in each cluster sharing a high similarity. 2) Cluster-oriented Constraint: During the training process, we first quantize the image patches using the rearranged codebook. For the output probability distribution $\hat{Y}$, we further compute the cluster-level distribution $\hat{Y}_{C}$ by applying LogSumExp operation for the probabilities in each cluster $\hat{Y}_{{\color[rgb]{0,0,0}jm}}\sim\hat{Y}_{{\color[rgb]{0,0,0}(j+1)m-1}}$. Then we compute the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ apart from the token-oriented cross-entropy loss $\mathcal{L}_{TCE}$, which ensures a high probability of the predicted token located in the correct cluster, thereby enhancing generation quality.
Observing that an inherent difference between LLM-based image generation and text generation lies in that text generation only requires a predicted text index that can be directly mapped into the corresponding word, while image generation essentially requires the image embedding that corresponds to the index, and then decodes it into an image, we further explore the characteristics of the image embedding space. We conduct a detailed analysis on the similarity of the image embeddings in the codebook learned by image tokenizers (Sec. 3), and find that similar embeddings convey similar information in the image space: When the embeddings of certain image patches are replaced by similar image embeddings, the decoded images are nearly identical. Therefore, we leverage this property to enhance the LLM-based image generation model.
We first propose a Codebook Rearrangement strategy that uses a balanced K-means clustering method to rearrange the embeddings in the codebook into clusters of equal sizes (Fig. 3 (left), Sec. 3.2), where the image embeddings in each cluster share a high similarity. Once the model predicts the correct cluster index (even if the token index is wrong), the corresponding embedding is located in the correct cluster and is thus similar to the target one, then the output images will be quite similar to the target image. To enable the model with the ability to predict the correct cluster index, we further propose a Cluster-oriented Cross-entropy Loss (Fig. 3 (right), Sec. 3.3). Our rearranged codebook and cluster-oriented cross-entropy loss ensure that, even if the model predicts incorrect tokens, there is a high probability of correctly predicting the cluster indices, which ensures high-quality generation results, largely improving the training efficiency and robustness of the LLM-based image generation model.
3.1 Analysis on Image Embedding Similarity
We first analyze the similarity of image embeddings in the codebook learned by image tokenizers, and find that similar embeddings (those that are close in distance) represent similar image information. I.e., when the embeddings of each image patch are replaced by other similar embeddings, the content of the decoded image remains almost unchanged. To verify this image embedding similarity property, we conduct detailed experiments on the codebook of VQGAN [14].
Specifically, an image $x$ is quantized into discrete tokens by: 1) using an image encoder $E$ (e.g., from VQGAN [14]) to extract features, obtaining $h\times w$ image features of dimension $C$, denoted as $\hat{z}=E(x)\in\mathcal{R}^{h\times w\times C}$; 2) quantizing each feature $\hat{z}^{(i,j)}$ into the code index $q^{(i,j)}$ of the nearest code in the codebook $\mathcal{Z}{\color[rgb]{0,0,0}\in\mathcal{R}^{N\times C}}$:
${\color[rgb]{0,0,0}q^{(i,j)}}=\arg\min\limits_{z_{k}\in\mathcal{Z}}\|\hat{z}^{(i,j)}-z_{k}\|{\color[rgb]{0,0,0}\in[0,N)},$ (1)
[…]
To measure the distance between two embeddings $z_{i}$, $z_{j}$ in the codebook, we define the “code distance” $D(z_{i},z_{j})$:
${\color[rgb]{0,0,0}D(z_{i},z_{j}):=d,\quad z_{j}=[\text{sort}(\{\|z_{k}-z_{i}\|\})]_{d},}$ (2)
calculated by measuring the Euclidean distance of each $z_{k}$ to $z_{i}$, sorting the distances, and $z_{j}$ is the $D(z_{i},z_{j})$-th closest embedding to $z_{i}$ among all embeddings in the codebook. A lower “code distance” indicates the two embeddings are closer in the embedding space.
Given an input image $x$, the quantized embedding $z_{q}$, and a generated image $\hat{x}^{\prime}=G(z_{q}^{\prime})$ decoded from $z_{q}^{\prime}$, we evaluate the similarity of the two images at different code distances $D(z_{q},z_{q}^{\prime})$. In Fig. 2 (a), we measure the image similarity by MSE distance and LPIPS [44], where lower scores indicate higher image similarity. Results show: 1) as the code distance increases, the image distance also increases; 2) but when the code distance is within a certain range (e.g., smaller than $12$), the image distance (MSE $0.104$ and LPIPS $66.44$) does not increase too much compared to the reconstructed image $G(z_{q})$ (whose code distance is 00 and has a $0.076$ MSE and $55.8$ LPIPS). We further visualize the images decoded from different code distances in Fig. 2(b), which shows the decoded images with lower code distances (e.g., smaller than $12$) look nearly identical to the source image and have a good visual quality. This indicates that even if the predicted token index is not the accurate target index, as long as the code distance between the corresponding embedding is within a certain range, the decoded image is similar to the desired one and shows a good visual quality. Therefore, we make use of this property to improve the generation quality and stability of the LLM-based visual generation model.
Algorithm 1 Balanced k-means Clustering
0:  Codebook $\mathcal{Z}$, number of clusters $n$
0:  The rearranged codebook $\hat{\mathcal{Z}}$ with local similarity
1:  Initialize:
2:   Compute the cluster size $m\leftarrow|\mathcal{Z}|/n$
3:   Randomly select $n$ codes as initial centroids $\{c_{j}\}_{j=1}^{n}$
4:   Initialize iteration counter $iter\leftarrow 0$
5:  repeat
6:     Compute the minimum distance $d_{i}$ from $z_{i}$ to $\{c_{j}\}_{j=1}^{n}$
7:     Sort $\{z_{i}\}$ in ascending order of $d_{i}$ to form $\mathcal{Z}^{\prime}$
8:     for each data point ${\color[rgb]{0,0,0}z_{i}}\in\mathcal{Z}^{\prime}$ do
9:        Compute the distance from $z_{i}$ to each centroid $c_{j}$
10:        Assign $z_{i}$ to the nearest cluster whose size does not exceed $m$
11:     end for
12:     for each cluster centroid $c_{j}$ do
13:        Update $c_{j}$ to the average of all embeddings assigned to cluster $j$
14:     end for
15:     Increment iteration counter $iter\leftarrow iter+1$
16:  until convergence or reaching $max\_iters$
17:  for each cluster centroid $c_{j}$ do
18:     Assign the embeddings in cluster $j$ to $\hat{\mathcal{Z}}_{[jm,(j+1)m)}$
19:  end for
20:  return  The rearranged codebook $\hat{\mathcal{Z}}$
3.2 Codebook Rearrangement
Rearrangement Target. In a codebook learned from the VQGAN [14] training process, the embeddings are often randomly distributed, with adjacent embeddings having no particular correlation, i.e., Euclidean distances between adjacent embeddings can be either close or far. Therefore, we first propose a Codebook Rearrangement strategy to reorder the codebook, so that the similarity between adjacent embeddings is maximized, ensuring that the surrounding embeddings of any given embedding exhibit high similarity.
Specifically, denote the codebook as $\mathcal{Z}=\{z_{i}\}_{i=1}^{N}$, which contains $N$ quantized embeddings $z_{i}$. To maximize the similarity between the adjacent embeddings, we aim to find a surjective mapping $M(\cdot)$ that satisfies:
$M=\arg\min\limits_{M}\sum_{i=1}^{N-1}{\color[rgb]{0,0,0}\|}z_{M(i)},z_{M(i+1)}{\color[rgb]{0,0,0}\|}.$ (3)
I.e., after reordering each embedding $z_{i}$ to index $M(i)$, the sum of distances between adjacent embeddings is minimized. And $\hat{\mathcal{Z}}=M(\mathcal{Z})$ is the rearranged codebook.
However, this optimization can be reduced to the Hamiltonian path problem (see #Suppl), and solving such a problem is NP-hard. Therefore, we try to relax this problem into a solvable form.
[…]
Balanced K-means Clustering. To cluster the embeddings in the codebook, we design a balanced K-means clustering method to uniformly divide the codebook into $n$ clusters, with the embeddings in each cluster close to each other.
As shown in Alg. 1, we first randomly select $n$ codes as the initial centroids $\{c_{j}\}_{j=1}^{n}$. Then, we iteratively update each cluster. In each iteration, 1) we first compute the minimum distance $d_{i}$ from each embedding $z_{i}$ to the $n$ centroids $\{c_{j}\}_{j=1}^{n}$. 2) We sort $\{z_{i}\}$ in ascending order of $d_{i}$ to form $\mathcal{Z}^{\prime}$ (i.e., embeddings closer to the nearest centroid has smaller index). This sorting and reordering step ensures that the embeddings closer to the nearest centroid are allocated first, and the farther ones are allocated later, which is necessary in cluster size balanced clustering. 3) Then, for each $z_{i}$ in the ordered codebook $\mathcal{Z}^{\prime}$, we compute its distance to each centroid $c_{j}$. 4) We assign $z_{i}$ to the nearest cluster whose size does not exceed $m$ (where $m=\frac{N}{n}$). 5) Finally, we update each cluster centroid $c_{j}$ to the average of the embeddings assigned to that cluster, and repeat the above process until it converges or reaches the pre-defined maximum iteration. After obtaining the final cluster centroids $c_{j}$ and $n$ clusters, where each cluster contains $m$ embeddings, we rearrange the codebook $\hat{\mathcal{Z}}$ by assigning the embeddings in cluster $j$ to indices $[jm,(j+1)m)$.
3.3 Cluster-oriented Visual Generation
Analysis on Token-oriented Cross-entropy Loss. In LLM-based visual generation models [38, 40], cross-entropy loss is the most commonly used loss in the training process. Denote the ground truth one-hot vector as $Y$ ($Y$ is a $N$-dimensional vector, with $Y_{i=y}=1$ and $Y_{i\neq y}=0$, where $y$ is the class label of the token), and the predicted probability distribution as $\hat{Y}{\color[rgb]{0,0,0}\in\mathcal{R}^{N}(\sum\hat{Y}_{i}=1)}$. The token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ is formulated as:
$\mathcal{L}_{TCE}=-\sum_{i=1}^{n}Y_{i}\log\hat{Y}_{i}.$ (4)
However, we observe that if a token index is wrongly predicted as a similar embedding’s index, this token-oriented cross-entropy loss will penalize the case, but the decoded image is actually not much different from the target image.
To tolerate this case, we leverage the rearranged codebook, where the embeddings in the same cluster are close to each other. In Sec3, it has been demonstrated that even if some image token indices are predicted incorrectly, as long as the embeddings of the predicted tokens are not far from the embeddings of the target tokens, the semantic integrity and quality of the generated image will not be significantly affected. Therefore, leveraging the reordered codebook from Sec3.2, a natural idea is to guide the model to first predict the cluster and then locate the specific token within that cluster. Since the number of clusters ($n$) is much smaller than the size of the codebook ($N$), predicting the cluster of a token is a much easier task than directly predicting a specific token. Once the cluster is accurately predicted, even if the specific token prediction is incorrect, the quality of the generated image can still be maintained.
Cluster-oriented Cross-entropy Loss. To guide the model to predict the correct cluster of the next token, we propose a cluster-wise cross-entropy loss $\mathcal{L}_{C}$. Specifically, we first derive the ground truth cluster label $y_{c}$ from the class label $y$ [^1] After codebook rearrangement, the class label has been updated accordingly. by $y_{c}=\lfloor\frac{y}{m}\rfloor$, where $m$ is the number of tokens in each cluster. Then, to form a new cluster-level probability distribution $\hat{Y}_{C}{\color[rgb]{0,0,0}\in\mathcal{R}^{n}(\sum\hat{Y}_{C,i}=1)}$ from the token-level probability distribution $\hat{Y}$, we employ the following operation to calculate the probability of each cluster, where within a cluster, the samples with larger probability contribute more to the cluster probability:
$\hat{Y}_{C,j}$ $=\nicefrac{{\sum\limits_{i=jm}^{(j+1)m-1}exp(\hat{Y}_{i})}}{{\sum\limits_{i=1}^{{\color[rgb]{0,0,0}N}}exp(\hat{Y}_{i})}},{\color[rgb]{0,0,0}\quad j=1,\ldots,n}$ (5)
where $\hat{Y}_{C,j}$ represents the probability that the generated token belongs to cluster $j$.
With the ground-truth cluster label $y_{c}$ and the predicted cluster-level probability distribution $\hat{Y}_{C,i}$, our cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ is formulated as:
$\mathcal{L}_{CCE}=-\sum_{{\color[rgb]{0,0,0}j}=1}^{n}Y_{C,{\color[rgb]{0,0,0}j}}\log\hat{Y}_{C,{\color[rgb]{0,0,0}j}},$ (6)
where $Y_{C}$ is the one-hot vector spanned by the label $y_{c}$. With our cluster-oriented loss, even if the model predicts the wrong token index, there is a high probability that the token is located in the target cluster, which ensures a high similarity between the output image and the target one, effectively improving the generation quality and robustness.
Final Loss Function. The final loss function $\mathcal{L}$ combines both the token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ and the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$:
$\mathcal{L}=\mathcal{L}_{TCE}+\lambda\mathcal{L}_{CCE}.$ (7)
4 Experiments
[…]
GAN BigGAN [5] 112M 6.95 224.5 0.89 0.38
GigaGAN [23] 569M 3.45 225.5 0.84 0.61
StyleGan-XL [37] 166M 2.30 265.1 0.78 0.53
Diffusion ADM [13] 554M 10.94 101.0 0.69 0.63
CDM [21] -- 4.88 158.7 -- --
LDM-4 [35] 400M 3.60 247.7 -- --
DiT-XL/2 [31] 675M 2.27 278.2 0.83 0.57
Mask. MaskGIT [7] 227M 6.18 182.1 0.80 0.51
MaskGIT-re [7] 227M 4.02 355.6 -- --
AR VQGAN [14] 227M 18.65 80.4 0.78 0.26
VQGAN [14] 1.4B 15.78 74.3 -- --
VQGAN-re [14] 1.4B 5.20 280.3 -- --
ViT-VQGAN [43] 1.7B 4.17 175.1 -- --
ViT-VQGAN-re [43] 1.7B 3.04 227.4 -- --
RQTran. [27] 3.8B 7.55 134.0 -- --
RQTran.-re [27] 3.8B 3.80 323.7 -- --
AR LlamaGen-B [38]  [^1] M 5.46 193.6 0.83 0.45
LlamaGen-L [38] 343M 3.29 227.8 0.82 0.53
LlamaGen-XL [38] 775M 2.63 244.1 0.81 0.58
LlamaGen-XXL [38] 1.4B 2.34 253.9 0.80 0.59
Ours IAR-B  [^1] M 5.14 202.0 0.85 0.45
IAR-L 343M 3.18 234.8 0.82 0.53
IAR-XL 775M 2.52 248.1 0.82 0.58
IAR-XXL (CFG=1.7) 1.4B 2.19 278.9 0.81 0.58
IAR-XXL (CFG=2.5) 1.4B 3.74 362.0 0.86 0.54
Table 1: Comparison between different types of image generation model on class-conditional ImageNet 256$\times$256 benchmark with FID, IS, precision, and recall.
4.1 Experimental settings
In this paper, we choose LLamaGen [38] as the main baseline and apply our method to it. We follow the official training details and keep the hyperparameters the same as the official ones. We conduct experiments on ImageNet [11]. To evaluate the model performance, we generate 50K images with randomly selected labels. Then, we compute the FID [18], IS [36], precision, and recall [26] for the generated data (see #Suppl for details).
Tokens Model #Para. 50 epoch 300 epoch
[…]
16 $\times$ 16 LlamaGen-B  [^1] M 7.22 178.3 0.86 0.38 5.46 193.6 0.84 0.46
LlamaGen-L 343M 4.20 200.0 0.82 0.51 3.80 248.3 0.83 0.52
LlamaGen-XL 775M 3.39 227.1 0.81 0.54 - - - -
LlamaGen-XXL 1.4B 3.09 253.6 0.83 0.53 - - - -
IAR-B  [^1] M 6.90 179.2 0.86 0.40 5.14 202.0 0.85 0.45
IAR-L 343M 4.10 207.1 0.82 0.51 3.40 271.3 0.84 0.51
IAR-XL 775M 3.36 228.9 0.82 0.54 - - - -
IAR-XXL 1.4B 3.01 257.4 0.83 0.53 - - - -
24 $\times$ 24 LlamaGen-B  [^1] M 8.31 154.7 0.84 0.38 6.09 182.5 0.84 0.42
LlamaGen-L 343M 4.61 191.4 0.82 0.50 3.29 227.8 0.82 0.53
LlamaGen-XL 775M 3.24 245.7 0.83 0.53 2.63 244.1 0.81 0.58
LlamaGen-XXL 1.4B 2.89 236.2 0.80 0.56 2.34 253.9 0.81 0.60
IAR-B  [^1] M 7.80 153.3 0.84 0.39 5.77 192.5 0.85 0.42
IAR-L 343M 4.35 197.2 0.81 0.51 3.18 234.8 0.82 0.53
IAR-XL 775M 3.15 228.8 0.81 0.54 2.52 248.1 0.82 0.58
IAR-XXL 1.4B 2.87 249.9 0.82 0.56 2.19 265.6 0.81 0.58
Table 2: Comparison with LlamaGen across different image tokens and model sizes. Following LlamaGen, we only train XL and XXL version on $16\times 16$ tokens for 50 epochs.
4.2 Comparison Experiments
Comparison with baselines. We compare our model with the GAN-based methods [5, 23, 37], diffusion-based methods [13, 21, 35, 31], mask-prediction-based methods [7] and autoregressive-based methods [14, 43, 27, 38] on ImageNet [11]. The results are shown in Tab. 2. Our method achieves the state-of-the-art FID (2.19) and IS (362.0). And for each model size, our IAR consistently outperforms LlamaGen, demonstrating the effectiveness of our cluster-oriented autoregressive visual generation strategy.
Figure 4: (a) Model performance (IAR-B and IAR-L) across different CFGs; (b) Model performance on different parameter numbers ( [^1] M to 3B) compared to LlamaGen; and (c) Model performance on different epochs compared to LlamaGen.
More comparisons with LlamaGen. We further compare our model with the baseline LlamaGen across different model sizes and image tokens to validate the effectiveness of our model. Specifically, we choose four models with model sizes ranging from  [^1] M to 1.4B parameters. And we also conduct experiments on $16\times 16$ image tokens (corresponds to $256\times 256$ image) and $24\times 24$ image tokens (corresponds to $384\times 384$ image). We choose the model trained with 50 epochs and 300 epochs for evaluation. The comparison results are shown in Tab. 2. It can be seen that across different model sizes, image tokens, and training epochs, our model can always improve the baseline with a better FID and IS, validating the effectiveness and robustness of our model in LLM-based visual generation.
Effects of classifier-free guidance (CFG). IAR can employ classifier-free guidance [19] to generate images with higher quality. We show the performance (FID & IS) of our model ( [^1] M and 343M parameters) on different CFGs ranging from 1.0 to 2.75 in Fig. 4 (a). It can be seen that compared to the model without CFG (i.e., CFG=1.0), the model with CFG can largely improve the FID, which means the model can generate images with much higher quality. Specifically, as the CFG increases, the FID first decreases and then increases, indicating that a too-large CFG may influence the generation quality, which is consistent with LlamaGen and previous diffusion-based methods [13, 19].
Effects of model size. We conduct experiments on different model sizes, where the parameter number ranges from  [^1] M to 1.4B. It can be seen in Tab. 2 and Fig. 4 (b) that as the parameter number grows, the model shows a clear improvement in the performance, which aligns with the scaling law [24], indicating that as the number of parameter increases, our IVR can be more effective.
Training efficiency. We compare the training speed between IAR-B and LlamaGen-B (for different model sizes, please refer to #Suppl). We evaluate the model performance on different epochs (from 50 to 300). The FID and IS results are shown in Fig. 4 (c), which shows that our 175-epoch IAR reaches almost the same FID score as the 300-epoch LlamaGen, indicating that our model can accelerate the training speed for almost $42\%$. And comparing the models trained on 300 epochs, our model effectively improves the generation quality, validating that our model can improve the training efficiency of the LLM-based visual generation model.
4.3 Ablation Study
To evaluate the effectiveness of our model, we conduct several ablation studies to validate each module separately. We conduct the ablation studies based on IAR-B ( [^1] M parameters), and train the model for 100 epochs on ImageNet.
Ablation on codebook rearrangement and $\mathcal{L}_{CCE}$. We first evaluate the effectiveness of the codebook rearrangement strategy and the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ We conduct experiments on three ablated models: the model without both codebook rearrangement and $\mathcal{L}_{CCE}$ (original LlamaGen), the model without codebook rearrangement, and the model without $\mathcal{L}_{CCE}$. The results are shown in Tab 3. It can be seen that the model with only the codebook rearrangement (w.o. CCE loss) performs almost the same as LLamaGen. This is because the original token-oriented cross-entropy loss does not consider cluster information and makes no distinction between embeddings with distant indices and those with close indices, thus failing to utilize the similarity within local clusters. Therefore, even with codebook rearrangement, the result remains almost the same as the original LlamaGen. For the model with CCE loss only (w.o. rearrangement), we treat adjacent embeddings in a window (e.g., $0\text{-}128$-th codes, $128\text{-}256$-th codes) as a “cluster”. Without the clustering and reorderng process, codes in such “clusters” are not similar to each other, therefore CCE loss cannot function well and only slightly improves the performance. In contrast, our model with both codebook rearrangement and $\mathcal{L}_{CCE}$ has the best performance in fitting the ImageNet dataset.
[…]
7.14 166.38 0.84 0.40
7.15 163.09 0.84 0.41
6.96 168.70 0.84 0.40
6.77 171.73 0.84 0.42
Table 3: Ablation studies on codebook rearrangement strategy and the cluster-oriented cross-entropy loss $\mathbf{\mathcal{L}_{CCE}}$.
Ablation on cluster size. A key hyperparameter in our model is the size of each cluster. A too-small cluster size may make it too hard for the model to predict the correct cluster index, which leads the model to perform similarly to the original LlamaGen (Note that the model with cluster size $1$ is the same as LlamaGen). In contrast, a too-big cluster size may cause the image embeddings in each cluster to be too different, such that even if the model predicts the correct cluster index, the gap to the target image embedding is still too big. Therefore, it is important to choose an appropriate cluster size to balance the inner similarity in each cluster and the ease of predicting the correct cluster index. We conduct experiments on different cluster sizes ranging from $1$ to $512$, where the codebook size is $16384$. We compute the inner MSE of cluster, i.e., the average distance from each code in a cluster to the centroid, and evaluate the model performance by FID, IS, Precision, and Recall in Tab. 4. It can be seen that as the cluster size increases, the inner MSE distance becomes higher, indicating that the image codes in each cluster become less similar. Moreover, the model reaches the best performance when the cluster size is $128$, which happens to be the square root of the codebook size.
[…]
1 0 7.14 166.38 0.84 0.40
8 0.018 7.02 167.09 0.85 0.40
16 0.022 6.99 166.39 0.84 0.42
32 0.028 6.83 170.10 0.84 0.40
64 0.034 6.92 171.65 0.85 0.39
128 0.041 6.77 171.73 0.84 0.42
256 0.050 6.81 170.58 0.85 0.41
512 0.059 6.94 163.54 0.83 0.42
Table 4: Ablation studies on different cluster sizes from $1$ to $512$, where cluster size $1$ corresponds to the official LlamaGen.
Ablation on weight of $\mathcal{L}_{CCE}$. Since we introduce an additional cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ during the training process, we conduct experiments to find the influence of its corresponding weight $\lambda$. We choose $\lambda$ ranging from $0.1$ to $1.5$ and evaluate the model performance in Tab. 5. It can be seen that a too-small weight will cause the model to perform similarly to the original LlamGen, while a too-big weight will make the model concern too much about the cluster index and overlook the ground-truth token index, which also makes the model perform worse. In summary, we find that when the weight $\mathbf{\lambda=1}$, the model can balance the two goals and achieves the best performance.
[…]
0 7.14 166.38 0.84 0.40
0.1 7.01 167.16 0.84 0.40
0.25 7.05 160.73 0.84 0.42
0.5 6.78 171.15 0.84 0.41
0.75 6.80 170.06 0.84 0.41
1 6.77 171.73 0.84 0.42
1.5 6.87 170.60 0.84 0.41
Table 5: Ablation studies on the weight $\lambda$ of the cluster-oriented cross-attention loss $\mathbf{\mathcal{L}_{CCE}}$.
5 Conclusion
In this paper, we analyzed the difference between the natural language and images in LLM-based visual generation. We find that similar image embeddings in the codebook can produce similar images. Based on this observation, we propose IAR, an Improved Autoregressive Visual Generation Method that effectively enhances the training efficiency and quality of the LLM-based visual generation model. We first conduct Codebook Rearrangement with a balanced K-means clustering algorithm to reorder the codebook into clusters with equal sizes, where the image embeddings within each cluster share similarities. Then, we introduce a Cluster-oriented Cross-entropy Loss to enable the model to learn the target cluster, which can guarantee a good generation quality even if the model predicts a wrong image token. We validate the effectiveness of our IAR on LlamaGen, and find that IAR can largely improve the training efficiency and generation quality across different parameter scales. Moreover, our IAR can be seamlessly applied to various LLM-based visual generation model, pointing out a new direction to improve the LLM-based visual generation field.
Appendix A Overview
The supplementary material is composed of:
  • Implementation details (Sec. B);
  • More details on optimization relaxation in codebook rearrangement (Sec. C)
  • Comparison with VAR [40] (Sec. D)
  • More analysis on our model (Sec. E);
  • More visualization results (Sec. F);
  • Future work (Sec. G).
We also provide the code of our model in the supplementary material, which can be easily applied to the existing autoregressive models.
Appendix B Implementation Details
Metrics. We employ four metrics to evaluate the effectiveness of the models:
  • Fréchet inception distance (FID) [18] measures the similarity between the features of the source data and the generated data according to their mean values and covariance. A smaller FID indicates better generation ability.
  • Inception Score (IS) [36] measures the quality and diversity of images by computing the information entropy of the generated images. A higher IS indicates better generation quality and diversity.
  • Precision/Recall [26] measures the class-conditional generation accuracy. A higher precision or recall indicates a better class-conditional generation performance.
Experiment settings. We follow the experiment settings as LlamaGen [38] and keep the hyperparameters consistent with it. The experiment details are shown in Tab. A1 and Tab. A2, where Tab. A2 is the inference settings for Tab. 2 of the main paper.
Sampling hyperparameters: Among the hyperparameters used in the inference process (Tab. A2), there are several important parameters, whose meanings are explained in detail below:
(1) Classifier-free guidance: Classifier-Free Guidance (CFG) [19] is originally a sampling method to improve diffusion models by combining conditional and unconditional score estimates. Beyond diffusion models, CFG can also be applied to the autoregressive image generation process [38]. Denoting the input image token sequence as $q$, our model as $\epsilon_{\theta}$, and the class condition as $c$, the autoregressive CFG is defined as:
$\tilde{\epsilon}_{\theta}(q,c)=(1+w)\epsilon_{\theta}(q,c)-w\epsilon_{\theta}(q,\phi),$ (A1)
where $\phi$ denotes the empty condition and $\epsilon(,)$ represents the predicted probability distribution for the next image token.
(2) Top-K: Top-K sampling [32] is a decoding strategy that selects tokens from the top $k$ highest-probability candidates. It focuses on the most likely tokens, but the fixed $k$ size may exclude important low-probability options.
(3) Top-P: Top-P sampling [22], also known as nucleus sampling, selects tokens dynamically from the smallest set whose cumulative probability meets or exceeds a threshold $p$. This approach adapts to the output distribution, balancing coherence and diversity in text generation.
(4) Temperature: In large language models (LLMs), temperature [2, 29] is a hyperparameter that controls the randomness of the generated token by adjusting the sharpness of the probability distribution: lower values make the output more deterministic, while higher values increase diversity and randomness. The probability $P_{i}$ for each token is calculated as:
$P_{i}=\frac{\exp{(l_{i}/T)}}{\sum_{j}\exp{(l_{j}/T)}},$
where $l_{i}$ represents the predicted probability distribution, and $T$ is the temperature.
Model B L XL XXL B L XL XXL
Parameter Num  [^1] M 343M 775M 1.4B  [^1] M 343M 775M 1.4B
Token Num 16$\times$16 24$\times$24
Optimizer AdamW
Weight decay 0.05
Learing Rate Scheduler Constant
Batch Size 256 256 256 256 256 256 256 512
Learning Rate 1E-04 1E-04 2E-04 2E-04 1E-04 1E-04 2E-04 2E-04
GPU Num 8 8 8 8 8 8 16 32
Epoch 300 300 50 50 300 300 300 300
FSDP No No No Yes No No No Yes
Table A1: The training settings and hyperparameters used in our model.
Model B L XL XXL B L XL XXL
Parameter Num 00000  [^1] M 343M 775M 1.4B  [^1] M 343M 775M 1.4B
Token Num 16$\times$16 24$\times$24
Batch Size 32
Random Seed 0
Top K 0
Top P 1.0
Temperature 1.0
CFG 2.0 2.0 1.75 2.0 2.25 1.75 1.75 1.65
Table A2: The inference settings and hyperparameters used in Tab. 2 of the main paper.
Classifier-free Guidance VAR-d16 + IAR VAR-d16
[…]
1.5 4.12 58.15 0.839 0.482 4.28 56.66 0.830 0.479
1.75 4.07 60.54 0.857 0.458 4.25 59.00 0.846 0.460
2.0 4.43 63.11 0.865 0.435 4.52 61.00 0.860 0.435
Table A3: Comparing VAR-d16 [40] with VAR+IAR on ImageNet [11]. It shows that our IAR also performs well in the next-scale prediction model, validating that our method can be widely applied to various autoregressive image generation models, enhancing their generative capabilities.
Classifier-free Guidance IAR-B IAR-L
[…]
1 29.70 43.96 0.566 0.632 20.56 62.96 0.595 0.666
1.5 10.69 103.59 0.732 0.532 4.39 178.78 0.778 0.566
1.75 7.43 135.55 0.783 0.501 3.18 234.79 0.824 0.530
2 6.06 165.22 0.822 0.454 3.49 279.09 0.855 0.499
2.25 5.77 192.45 0.850 0.421 4.43 311.08 0.873 0.466
2.5 6.11 213.76 0.869 0.381 5.61 340.18 0.890 0.425
2.75 6.73 232.35 0.884 0.360 6.74 358.48 0.898 0.401
Table A4: The Quantitative metrics on our model under different classifier-free guidance scales.
Model Size Epoch LlamaGen IAR
[…]
B Version 50 8.67 136.62 0.818 0.413 7.80 153.31 0.839 0.394
100 7.26 152.50 0.827 0.416 6.77 171.73 0.839 0.416
200 6.54 167.82 0.833 0.428 5.86 185.28 0.845 0.428
300 6.09 182.54 0.845 0.416 5.77 192.45 0.850 0.421
L Version 50 4.25 191.46 0.819 0.504 4.35 197.23 0.819 0.507
100 3.96 199.96 0.803 0.532 3.81 205.63 0.805 0.528
200 3.33 219.57 0.804 0.538 3.31 225.95 0.814 0.551
300 3.29 227.83 0.818 0.532 3.18 234.79 0.824 0.530
Table A5: The Quantitative metrics on our model and LlamaGen under different training epochs. The B version employs CFG=2.25 and the L-version employs CFG=1.75.
Figure A1: The training loss curves for the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ (a) and token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ (b) on $\mathbf{24\times 24}$ image tokens.
Figure A2: The training loss curves for the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ (a) and token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ (b) on $\mathbf{16\times 16}$ image tokens.
Appendix C Complexity Analysis of Codebook Rearrangement Target
In Sec. 3.2 of the main paper, we aim to rearrange the codebook such that the neighboring embeddings are as close to each other. We summarize this code rearrangement problem as an optimization problem, where we aim to find a surjective mapping $M(\cdot)$ that satisfies:
$M=\arg\min\limits_{M}\sum_{i={\color[rgb]{0,0,0}1}}^{N-1}\|z_{M(i)},z_{M(i+1)}\|.$ (A2)
After reordering each embedding $z_{i}$ to index $M(i)$, the sum of distances between adjacent embeddings is minimized. And $\hat{\mathcal{Z}}=M(\mathcal{Z})$ is the rearranged codebook.
However, this optimization can be reduced to the Shortest Hamiltonian path problem, which is a classical NP-hard problem. The Shortest Hamiltonian Path Problem is a variation of the Hamiltonian Path Problem. Its goal is to find a path that visits each vertex exactly once and minimizes the total weight (or distance) of the path. Formally, given a weighted graph $G=(V,E)$ with a weight function $w:E\rightarrow\mathbb{R}^{+}$, the goal is to find a Hamiltonian path $\pi^{*}=(\pi_{1},\pi_{2},\ldots,\pi_{N})$ such that the sum of the weights of the edges in the path, i.e., $\sum_{i=1}^{N-1}w(\pi_{i},\pi_{i+1})$, is minimized, which is formulated as:
$\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{N-1}w(\pi_{i},\pi_{i+1}).$ (A3)
Next, we prove that solving the optimization problem in Eq. (A2) can be reduced to the Shortest Hamiltonian path problem in Eq. (A3):
Proposition: Solving Eq. (A3) $\leq_{p}$ Solving Eq. (A2)
Proof.
Step 1: Construct a Complete Weighted Graph:
Define a complete graph $G=(V,E)$ where the vertex set $V=\{0,1,\ldots,N-1\}$ corresponds to the $N$ embeddings in the codebook $\mathcal{Z}$. Each edge $(i,j)\in E$ is assigned a weight $w(i,j)$ that equals to the distance between embeddings $z_{i}$ and $z_{j}$:
$w(i,j)=\|z_{i},z_{j}\|.$ (A4)
Step 2: Find the Minimum Weight Hamiltonian Path:
Finding the shortest Hamiltonian path $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{N})$ in $G$ aims to minimize the total weight:
$\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{N-1}w(\pi_{i},\pi_{i+1}).$ (A5)
Step 3: Mapping to the Original Problem:
The shortest Hamiltonian path $\pi^{*}$ provides the optimal permutation $M^{*}$ for the optimization problem in Eq. (A2), where $M^{*}(i)=\pi_{i}$, for $i=1,2,\ldots,N$.
In summary, the original optimization problem (in Eq. (A2)) of finding the optimal surjective mapping $M(\cdot)$ to minimize the sum of distances between consecutive embeddings, can be reduced to finding the minimum weight Hamiltonian path in a complete weighted graph, where the weights are given by the distances between embeddings.
Therefore, the original optimization problem is also NP-hard. And it is necessary to relax this optimization target to a clustering problem (main paper Sec. 3.2), which ensures the embeddings with a cluster share high similarities.
Appendix D Comparing IAR+VAR with VAR
VAR [40] extends the next-token prediction in autoregressive image generation to next-scale prediction, enabling the model to generate images progressively from small to large scales. At each scale, VAR predicts all tokens simultaneously, significantly enhancing the inference speed of the autoregressive image generation process. Our design is independent of the model structure in autoregressive image generation, allowing us to integrate our IAR with VAR, referred to as VAR+IAR. Given that most official VAR models are trained on 256 A100 GPUs, which is highly resource-intensive, we only train the VAR-d16 model for 100 epochs on ImageNet [11] and subsequently compare it with VAR+IAR.
Both models (VAR and VAR+IAR) are trained for 100 epochs with a batch size of 768, maintaining the same hyperparameters as the official VAR code. We then evaluate the trained models on different CFGs. The results, presented in Tab. A3, demonstrate that incorporating IAR into VAR enhances the original VAR in terms of generation quality and diversity, as evidenced by improved FID and IS scores. This validates the effectiveness of our model across different autoregressive image generation frameworks, showing the great potential of our IAR in the field of autoregressive image generation.
Appendix E More Analysis on Our Model
Comprehensive metrics for models under different CFGs. This section presents the comprehensive metrics (FID, IS, Precision, Recall) for the models compared in Fig. 4 (a) of the main paper. As shown in Table A4, an increase in CFG leads to higher IS and precision, while recall decreases. Unlike these three metrics, FID initially improves and then deteriorates, achieving its optimal value at an intermediate CFG. Furthermore, the optimal CFG for FID varies with model size (e.g., CFG=2.25 for IAR-B and CFG=1.75 for IAR-L).
Comprehensive metrics for models under different training epochs. Table A5 presents a comparison between our IAR and LlamaGen [38] over various training epochs, illustrating that our model consistently outperforms LlamaGen at all stages of training. Notably, the 200-epoch IAR-B exceeds the performance of the 300-epoch LlamaGen-B, while the 200-epoch IAR-L performs similarly to the 300-epoch LlamaGen-L, highlighting the high training efficiency of our model. (Note that all B-version models use CFG=2.25, whereas all L-version models use CFG=1.75)
Training losses for different model sizes. We show the training loss curves for both the two losses: 1) cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ and the token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ when training on $24\times 24$ image tokens (Fig. A1) and $16\times 16$ image tokens (Fig. A2). It can be seen that as the model size increases, both the two losses decrease faster and converge to a lower value, which aligns with the scaling law [24]. Note that since we follow the training setting of LlamaGen [38], we only train IAR-XL and IAR-XXL on $16\times 16$ image tokens for 50 epochs.
Effectiveness of $\mathbf{\mathcal{L}_{CCE}}$. In the main paper, we introduce the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$, designed to enhance the model’s awareness of cluster information, thereby increasing the likelihood of predicting tokens within the target cluster. It is hard to directly illustrate the effectiveness of $\mathcal{L}_{CCE}$ by its loss value directly. Therefore, we design an alternative way where we visualize the loss curves for token-oriented cross-entropy loss $\mathcal{L}_{TCE}$ and their corresponding FIDs for LlamaGen-B and our model in Fig. A3. The results indicate that, compared to LlamaGen, our model exhibits a higher token-oriented cross-entropy loss but achieves a superior FID. This suggests that our model has slightly lower token-oriented prediction accuracy, which is expected since the introduction of $\mathcal{L}_{CCE}$ partially diverts the original loss $\mathcal{L}_{TCE}$. Therefore, the improvement of FID can only come from our proposed cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$. Since $\mathcal{L}_{CCE}$ effectively increases the likelihood of predicting the correct cluster, combined with the embedding similarities within the cluster, it ultimately leads to the generation of images with better FID, demonstrating the efficacy of $\mathcal{L}_{CCE}$ in our model.
Figure A3: Comparison between LlamaGen-B and ours on the token-oriented cross-entropy loss $\mathbf{\mathcal{L}_{TCE}}$ and the FID score in different training iterations. Our model has a higher $\mathcal{L}_{TCE}$ than that of LlamaGen but achieves a better FID, indicating the effectiveness of our cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$.
00000000000Model IAR LlamaGen [38]
B L XL XXL B L XL XXL
Cluster-level Accuracy (Top-1, %) 15.54 17.12 18.01 19.02 13.44 14.81 15.71 16.49
Cluster-level Accuracy (Top-5, %) 41.48 44.68 46.30 48.29 30.37 33.13 34.87 36.38
Token-level Accuracy (Top-1, %) 2.62 3.17 3.56 3.88 2.64 3.19 3.59 3.95
Token-level Accuracy (Top-5, %) 7.34 8.86 9.90 10.75 7.37 8.91 9.98 10.96
Table A6: Comparison of the token-level prediction accuracy, cluster-level prediction accuracy, and the embedding-level MSE distance between our IAR and LlamaGen.
Token prediction accuracy. We compute the token prediction accuracy of our model and LlamaGen [38] on different model sizes ($24\times 24$ tokens). Specifically, for an image token sequence $q=\{q^{1},q^{2},\cdots q^{576}\}$ with corresponding image embedding sequence $z_{q}=\{z_{q}^{1},z_{q}^{2},\cdots z_{q}^{576}\}$ and class label $c$, we enumerate $i$ from $1$ to $575$ and predict $\hat{q}^{i+1}\sim P^{i+1}=\epsilon_{\theta}(\hat{q}^{i+1}|c,q^{1},q^{2},\cdots q^{i})$ using the model $\epsilon_{\theta}$. We then compute the Top-1 and Top-5 accuracy $Acc^{i}$ between $\hat{q}^{i+1}$ and the ground truth $q^{i+1}$. The average accuracy for an image is calculated as $Acc=\frac{1}{575}\sum_{i=1}^{575}Acc^{i}$. Finally, we compute the cluster-level accuracy and token-level accuracy for all images in ImageNet [11] and record the average values in Tab. A6. Specifically, to compute the cluster-level accuracy for LlamaGen, we employ the balanced K-means clustering algorithm to decompose the codebook into $n$ clusters and then determine the target cluster index. We then assess whether the predicted token is located in the target cluster, thereby obtaining the cluster-level accuracy. From Tab. A6, it can be seen that our cluster-level accuracy is higher than that of LlamaGen, indicating the effectiveness of our cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$. Although our token-level accuracy is slightly lower, this is expected as the newly included loss $\mathcal{L}_{CCE}$ affects the original token-oriented cross-entropy loss $\mathcal{L}_{TCE}$, resulting in a slight decrease in token-level accuracy. However, our model still achieves better FID and IS compared to LlamaGen, further validating the effectiveness of our cluster-oriented token prediction strategy.
Appendix F More Visualization Results
[…]
Appendix G Future Work
The main idea of our IAR is to ensure a high similarity between the predicted image embedding and the target embedding, so that even if the model incorrectly predicts the target token, the output image still closely resembles the target image. This can be naturally considered as a continuous constraint on the image embedding, aiming to minimize the distance between the predicted and target image embeddings. However, this approach cannot be easily applied to LLM-based image generation models due to the non-differentiable nature of the embedding quantization operation. Therefore, this paper relaxes the problem into a cluster-oriented token prediction problem, which can be easily integrated into the current autoregressive image generation model. We believe that in future work, employing this continuous constraint in autoregressive image generation may further enhance the model performance.
Figure A4: The generated images for alp, promontory, volcano, and coral reef by IAR-XXL with 4.0 CFG.
Figure A5: The generated images for sports car, balloon, convertible, and space shuttle by IAR-XXL with 4.0 CFG.
Figure A6: The generated images for castle, church, beacon, and cinema by IAR-XXL with 4.0 CFG.
Figure A7: The generated images for bridge, ocean liner, white stork, and Pomeranian by IAR-XXL with 4.0 CFG.